OpenStudy (anonymous):

Center(-4,-11) and radius 7 units

OpenStudy (anonymous):

Write the equation for a circle that satisfies the given conditions.

OpenStudy (anonymous):

the formula for a circle is (x-x1)^2+(y-y1)^2 = r^2 where (x1, y1) is the ordered pair of the center of the circle, and r is the radius

OpenStudy (anonymous):

SO ITS \[\left( X+4 \right)^{2}+\left( Y+11 \right)^{2}=49\]

OpenStudy (anonymous):

yes, that looks correct!

OpenStudy (anonymous):

COOL THANKS FOR UR HELP

OpenStudy (anonymous):

yw